OpenStudy (anonymous):

how much 5% acid must be mixed with 15% acid to obtain 800ml of 7% acid

OpenStudy (amistre64):

x+y = 800; y = 800-x x(.05) + y(.15) = 800(.07)

OpenStudy (amistre64):

or x = 800-y :)

OpenStudy (amistre64):

40 -.05y + .15y = 56 .10y = 16 y = 160; 160ml (15%) x = 800-160 = 640(5%)